OpenStudy (cathyangs):

If you toss two six-sided die, what is the probability of getting at least one 3?

OpenStudy (anonymous):

11/36

OpenStudy (anonymous):

5/6*1/6*2+1/6*1/6

OpenStudy (cathyangs):

what does that work out to? I can't really understand the way you typed it!

OpenStudy (anonymous):

it works out to 11/36

OpenStudy (anonymous):

\[\frac{5}{6}\cdot\frac{1}{6}\cdot2+\frac{1}{6}\cdot\frac{1}{6}\]

OpenStudy (cathyangs):

ah... Thank you... How does that work though?

OpenStudy (anonymous):

you can get one three OR two three, so: first possibility: first throw: 3; probability 1/6 second throw: anything except 3; probability 5/6 total first probability 1/6*5/6 second possibility: first throw: anything except 3; probability 5/6 second throw: 3; probability 1/6 total second probability 5/6*1/6 third possibility first throw: 3; probability 1/6 second throw: 3; probability 1/6 total third probability 1/6*1/6 and now add everything

OpenStudy (anonymous):

You get 1, 3 2, 3 ... 6, 3 6 choices 3, 1 3, 2 3, 4 3, 5 3, 6 3,3 from this list is excluded because there's a same one from the previous list. 5 choices 6 + 5 = 11

OpenStudy (cathyangs):

but aren't you double counting? The order doesn't matter in this situation, right?

OpenStudy (anonymous):

I didn't double count. no, the order matters since there are two dices

OpenStudy (anonymous):

it matters, because it's different

OpenStudy (cathyangs):

but wouldn't 1,3 be the same thing as 3, 1, because you toss the die at the same time?!? I don't know the answer to this question, and I won't be told until January, so THANK YOU for helping me! :D

OpenStudy (anonymous):

probability is equal but it's not the same

OpenStudy (anonymous):

This is a basic probability question

OpenStudy (anonymous):

you can try simple experiment with tossing coins and checking how many times you get at least 1 head, so 1head and 1 tail or 1 tail and 1 head or 1 head and 1 head

OpenStudy (anonymous):

you will see that probability will be somewhere near answer

OpenStudy (cathyangs):

Thanks for that help... Then this is a permutation, not a combination? If they are not the same, then order does matter?

OpenStudy (anonymous):

with permutations and combination you don't calculate probabilities

OpenStudy (cathyangs):

I'd rather not say. So to find the total number of solutions with 3s in them...I think I get your way(s), though my friend says that it's: 5 ways to get one 3, and 1 way to get two 3s. this is 6 over: (6*6)/2 which is 6/18, or 1/3. I am slightly confuzzled.